A fund’s reported average return is 10% per year and it charges a 5% front-end load. An investor has $50,000 to invest with the fund for a 5-year horizon. What is the investor’s average return over the 5 years?

8.87%

10.6%

9%

27.62%
Answer #1

Gross Amount for investment = $50000

Front load = 5%* $50000 = $2500

Net Amount left for investment = $50000 - $2500 = $47500

Returns @ 10% = P*(1 + r)^t = $47500*(1+10%)^5 = $76499.23

Average return = (76499.23/50000)^(1/5) - 1 = 8.87% [Ans-(a)]